OpenStudy (anonymous):

need help 49x = 7x – 3

OpenStudy (anonymous):

subtract 7x from both sides 49x-7x=7x-7x-3 42x=-3 divide both sides by 42 42/42*x=-3/42 x=-1/14

OpenStudy (anonymous):

im sorry its suppose to be exponents 

OpenStudy (misty1212):

HI!!